# Sharp Tools: How Developers Wield Agentic AI in Real Software Engineering Tasks

### v2

@@ III-E2 Pre-Study Questionnaire
This questionnaire consisted of optional questions about participant demographics, as well as prior experience with AI tools. Prior experience questions were based on questions from previous studies of human-AI interaction [21] and developer-AI interaction [26]. We did this to explore how prior experience with AI tools might be related to participants’ use of the tool [26]. We based the list of the AI tools on those reported in the 2024 Stack Overflow Developer Survey [1]. We also included a question about participants’ perceptions of AI tools for programming in terms of usability [27], by adapting the Technology Acceptance Model questionnaire [45].
@@ III-F Analysis Approach
Based on relevance of actions to the immediate state of the chat with the agent, we interpret the presence and ordering of codes on participant actions at the prompt level (e.g., verifying code changes, stopping the agent) or at the issue level (e.g., writing code, running tests) and analyze codes on chat interactions at the prompt level. Since participants were instructed to think aloud, we do not analyze the durations of their activities, nor do we analyze the absolute counts of these activities to avoid biases toward participants who performed actions repetitively in shorter intervals. We also note specific utterances and scenarios that indicate communication barriers, reporting only those that recur across different participants.
While our results are largely qualitative, we include some numerical data throughout to give the reader a fuller sense of the trends that we observed. However, given the exploratory nature of the study and the relatively small sample sizes, we caution the reader against broad extrapolation or generalization. Congruently, we present findings descriptively, without applying inferential statistics.
Towards helping to ensure the validity of our findings, we sent an early version of our results to all participants as a member check [46]. Our study protocol has undergone privacy assessment and review by our organization. To aid in replication, qualitative codebooks and study materials can be accessed in our Supplementary Material.
@@ III-G Limitations
Another limitation is small sample size (19 participants); we accepted this limitation so that we could perform time-consuming, thorough action and trajectory coding, which does not easily scale up to more participants.
Most participants in our study had not previously used Cursor or any other SWE agent, and thus our results mostly reflect early usage, rather than regular long-term use. Our study is also subject to observer effects; as participants knew that we were studying their interactions, they might may have used the tool differently than in a private setting. Likewise, our instruction that participants try to use AI for all their tasks means that, in the wild, developers would likely be more prone to abandonment and shifting to manual effort. Our choices of Cursor in early 2025 also has an effect on generalizability; since AI tools and models are constantly evolving, some findings reflect present-day limitations of agents and their underlying models.
@@ IV Results
Figure 3 visualizes participants’ success and time spent as they submitted 269 prompts to the agent across 33 issues. Participants were satisfied with the final patch for 16 of these tasks, while 4 of these tasks were abandoned for external reasons (such as participants being unconvinced of the issue’s validity) for a total success rate of 55% $(16/(33-4))$. In this section, we first describe the patterns we observed when participants collaborated with the agent, followed by the communication challenges they faced in this collaboration. We then discuss how these patterns and other factors were associated with participants’ success.
@@ IV-A RQ1: How developers collaborate with SWE agents
In this subsection, we describe the strategies participants applied when collaborating with the agent and the patterns this led to, ranging from high-level patterns related to delegation and iteration to lower-level patterns related to the type of information participants provided and the outputs they sought from the agent.
@@ IV-A1 Delegation Strategies
  • If the agent is unable to generate a comprehensive fix or verify its changes, participants must then both manually understand the generated fix and then iterate on it to get to a valid fix. We observed that this iteration can be time-consuming, as the it can entail not only understanding the code change itself, but also understanding the reasoning behind the change, navigating between code changes, and understanding changes to test files.
In the other approach (the incremental resolution strategy), participants manually divide the task into sequential sub-tasks and ask the agent to solve each sub-task sequentially in separate prompts. When the participant is satisfied with the solution to one sub-task, they then ask the agent to work on the next sub-task. Although this approach is safer than the one-shot strategy because mistakes made by the agent can be reviewed and caught earlier, it requires more proactive involvement and communication from the participant to divide the issue into these sub-tasks. As evidence, participants who used this strategy used an average of 11.0 prompts per issue, compared to 7.0 for the one-shot strategy. Further, these participants were more likely to manually read existing code to enhance their own understanding while working on issues, doing so in 83% (15 out of 18) of the issues they worked on as compared to 60% (9 out of 15) for other participants.
One participant (P7), who had used Cursor before, explained that they preferred different strategies in different scenarios: “For an issue that is very contained, I trust it to come up with the changes itself, but for issues that are more involved with different parts of the codebase I feel have to point it in the right direction so that anything sensible comes out at the end”. However, we also observed that these approaches are not mutually exclusive; when delegating the entire issue to the agent via the one-shot strategy, an incomplete solution may require the participant to prompt the agent with extra steps via incremental resolution. P3 noted how they dynamically changed their level of involvement during the collaboration after starting with the one-shot strategy: “The AI agent was able to generate good structure… once more context was needed, it committed errors that need more handholding and additional time to be spent on and guide it”.
@@ IV-A2 Knowledge Transfer
When working with the agent to solve issues, we observed that participants provided two different types of information to the agent to help it make the desired code changes:
@@
  • Contextual knowledge from the issue description and the environmental context (such as test logs). Eg: “There are build errors in your solution”
@@
  • Expert knowledge relating to code implementation or convention that cannot be observed directly from context or the issue description and is based on the developer’s prior knowledge of the repository. Eg: “Usually we don’t set an upperbound. I think 3.10 should be fine now”
@@
Participants’ patterns of providing such knowledge to the agent mirror those of software engineering managers. Kalliamvakou and colleagues [47] report that when assigning work, great managers in software engineering leave the implementation details to the engineer, providing actionable feedback when required. Analogously, participants tended to provide less expert knowledge to the agent up-front, and more as feedback; participants provided expert knowledge in 49% of prompts that asked the agent to generate a new code change, compared to in 65% of prompts that ask the agent to refine a code change (i.e., iterate upon a previously generated change). This relationship was even stronger for the 12 participants in managerial roles – they provided expert knowledge in 45% of their prompts seeking new changes versus 75% seeking to refine changes. As shown in Figure 4, the analogy is reflected in the post-study questionnaire, in which participants report viewing the agent more as an assistant (at a lower level in the organizational hierarchy) than as a colleague (at the same level) or advisor (at a higher level).
@@
While participants usually tended to provided expert knowledge in feedback prompts rather than up-front, we found that experience factors moderated this relationship. Participants who had more familiarity with the repository tended to provide expert knowledge more often – participants with over 10 commits to the repository in the four months leading up to the study provided such insights 70% of time when seeking any code changes as compared to 48% for other participants. Similarly, participants who had previous experience with Cursor also provided more expert knowledge, doing so in 81% of prompts seeking code changes. These participants also tended to provide expert knowledge more often as feedback – they did so for 89% of prompts trying to refine a code change as compared to for 67% of prompts when asking for a new code change. Further, participants applying the one-shot strategy provided expert knowledge less often (in 47% of their prompts seeking code changes) as compared to those applying the incremental resolution strategy (did so in 64% of their prompts seeking code changes).
IV-A3 Tasks Requested
@@
We observed that participants not only requested the SWE agent to make code changes to resolve the the entire task specified in the GitHub issue but also requested assistance in sub-tasks corresponding to other activities of the software development process. While 50% of participants’ prompts requested code changes, 27% of the prompts sought explanations of details related to the existing codebase (code comprehension), 16% of these prompts sought to run tests and 11% of these prompts sought explanations related to changes made by the agent (code review). Note that the above categories are not mutually exclusive. This distribution of prompts reflects participants’ perceptions of the agent as reported in the post-study questionnaire (Figure 4). Participants viewed the agent the most as a content generator; prompts requesting code generation were most common. Since ‘reference guide’ was a close second, participants also often used prompts to ask the agent about the codebase. On the other hand, participants did not view the agent as a reviewer, and were thus less likely to ask it to run tests or review its changes.
@@ IV-A4 Manual Actions
Participants’ perceptions of the agent shaped not only the types of prompts they provided the agent, but also the role the participants themselves took in the collaboration, tending to work on verification (debugging and testing) more than on content generation (writing code). Since participants reported not thinking of the agent as a code reviewer (Figure 4), they often took the responsibility to manually debug and test code upon themselves, doing so in 21 out of 33 issues by running tests or examining execution logs. In the case of bugs concerning UI changes, it was also easier for the participant to test changes manually by simply observing the current state of the UI rather than uploading images to the chat for the agent to interpret.
On the other hand, since the participants tended to view the agent more as a content generator, they were more likely to let the agent handle all the code changes to be made, manually writing code in only 14 issues. Further, in 10 of these issues, participants’ code interventions were limited to editing code suggested by the agent rather than adding new functionality.
Some participants expressed desire for AI tools to provide more support in reviewing and verification actions such as localizing, debugging, and testing, since these aspects of the software development process were often critical to resolving the issue. P12 expressed that that the agent was not particularly helpful in localizing the issue, mentioning that “I think the main part [of solving the issue] was finding where to put the code, so it might have been marginally faster to do it myself”. Similarly, participants mentioned the challenges they faced when collaborating with the agent for debugging; for instance, P8 responded in the post-study questionnaire that the agent was “not too great or useful at the debugging and converting-a-repro-into-a-test loop”, and P11 responded that “It’s puzzling to me why [the agent is] so good at generating code but bad at dealing with basic environmental commands”. Participants’ manual actions thus reflected not only their perceptions of the agent, but also a forced response to the constrained capabilities of the agent as a collaborator.
IV-A5 Reviewing Outputs
The agent provided output to communicate with participants in three ways:
  • Execution Trajectory: a trace of the agent’s work that appears while the agent is working, including context gathering steps such as codebase searches and files read (Figure  1b), along with actions taken such as terminal commands executed and files edited. (the bulk of Figure 1).
@@ IV-A5 Reviewing Outputs
We observed that participants tended to first track the agent’s execution live as-it-happened before reviewing explanations and diffs, with a preference for reading diffs. While Cursor enables its users to do manual tasks in parallel to its own execution, participants instead followed the agent’s execution as it happened in real time for 84% of their prompts. In some cases (23% of prompts), this was sufficient validation for participants, as they did not further verify the diffs or agent explanations. Usually, this live analysis was at least sufficient for participants to understand the agent’s context gathering steps; after following the agent’s execution live, they only reviewed this context gathering process for 7% of the agent’s responses. However, in the majority of cases (specifically, after 75% of the agent’s responses), participants reviewed diffs or explanations after following its execution in real time.
When reviewing code outputs, participants preferred directly verifying the generated code rather than reviewing code explanations. While participants followed the agent’s execution in real time for 95% of the responses that included code changes, these code changes often required further verification – participants reviewed diffs after 67% of such agent responses, while they only reviewed explanations after 31% of these responses. Of these 31%, participants verified both the diff and the explanations after 23% of such responses. This preference to review code is also reflected in the post-study questionnaire, in which participants ranked the ability to see detailed explanations of code changes as the least important feature among the different features of the agent (Figure 5). Such manual verification was usually sufficient for participants to have clarity on the agent’s code changes – they only asked for explanations of generated changes in 16% of their follow-up prompts.
@@
IV-A6 Handling Incorrect Outputs
To address incorrect code generated by the agent, participants in our study preferred to refine the agent’s changes by iterating on these changes rather than abandoning them and starting over. Such refinement was required more often when working on feature requests – participants used an average of 11.2 prompts when working on feature requests as compared to 6.3 when working on bugs. Overall, participants used an average of 8.2 prompts per issue, out of which 52% of prompts that sought code changes sought refinements to previous changes by the agent. Thus, the agent’s first outputs were often unsatisfactory and required further modifications. Despite this, participants only rejected code changes 10% of the time after the agent generated code changes. Further, while the agent allowed for participants to reset the chat to earlier points in the conversation (Figure 1a) – a feature we mentioned in the warm-up task – participants only used this feature in 15% of the issues they worked on. Additionally, participants only terminated the agent’s execution before completion (Figure 1e) in 11% of prompts. Thus, participants did not tend to backtrack by stopping or resetting the agent, instead preferring to iterate on imperfect changes.
Despite participants’ proclivity to iterate on incorrect outputs, some participants indicated that they did not think that iterating was always a good strategy. P14 mentioned of the agent in their post-study questionnaire that “if you continue to ask it to fix issues that it hallucinated on you might end up wasting time following in on the wrong path. It would be better to just explain the task in more detail or revert back to a previous breakpoint.” Similarly, P6 noted that the agent “was overly generative, not understanding the implications of previously written bad code, and doubling down by adding more code”.
IV-B RQ2: Developer-Agent Communication Barriers
@@ IV-B1 Lack of tacit knowledge
The agent was sometimes an ineffective collaborator because it lacked tacit knowledge, a type of knowledge that is gained through personal experience, rather than through explicit documentation. Rus and Lindvall [5] show that tacit knowledge comprises the majority of software engineering knowledge. The lack of such knowledge can make it difficult for new collaborators (such as the agent) to work effectively as they lack the knowledge that developers who are experienced with the repository, such as the participant, may have. Further, unlike expert insights into implementation details, it is difficult for the agent to learn tacit knowledge, as it is difficult to know when and how to communicate such information.
For example, in the case of P7 (who had prior experience with Cursor), the participant was able to interpret a temporal cue but unable to communicate their finding to the agent. When a terminal command running unit tests that was started by the agent seemed to be taking a long time to finish execution, the participant realized based on their experience with the repository that the tests were failing. However, rather than stopping the agent’s execution and explaining this in a prompt, they chose to work on the issue manually while the agent came to this realization by itself. In other cases, the agent suffered from a lack of social context. For example, the agent once offered an explanation of the codebase that seemed to contradict comments on the issue by a maintainer of the repository, making participant P8 very skeptical of the agent’s correctness. Prior knowledge of this context might have led the agent to question its assumptions before generating such an explanation.

### v3

@@ III-E2 Pre-Study Questionnaire
This questionnaire consisted of optional questions about participant demographics and prior experience with AI tools. We asked questions on prior experience to explore how this may be related to participants’ use of the tool [26]. These questions were based on previous studies of developer-AI interaction [21, 26]. We also included a question about participants’ perceptions of AI tools for programming in terms of usability [27] by adapting the Technology Acceptance Model questionnaire [53].
@@ III-F Analysis Approach
Based on the relevance of actions to the immediate state of the chat with the agent, we interpret the presence and ordering of participant actions at prompt-level granularity (e.g., verifying code changes, stopping the agent) or issue-level granularity (e.g., writing code, running tests) and analyze codes on chat interactions at the prompt-level granularity. Since participants were instructed to think aloud and study sessions varied in duration, we do not analyze the durations of participants’ activities, nor do we analyze the absolute counts of these activities to avoid biases toward participants who performed actions repetitively in shorter intervals. We also note specific utterances and scenarios that indicate barriers to effective collaborations, reporting only those that recur across different participants.
While our results are largely qualitative, we include some numerical data throughout to give the reader a fuller sense of the trends that we observed. However, given the exploratory nature of the study and the relatively small sample size, we caution the reader against broad extrapolation or generalization. Congruently, we present findings descriptively, without applying inferential statistics.
Towards helping to ensure the validity of our findings, we sent an early version of our results to all participants as a member check [54]. Our study protocol has undergone privacy assessment and review by our organization. To aid in replication, qualitative codebooks and study materials can be accessed in the Appendix.
@@ III-G Limitations
Another limitation is small sample size (19 participants); we accepted this limitation so that we could perform time-consuming, thorough action and trajectory coding which does not easily scale up to more participants. However, since participants in our study worked on open issues that did not have predefined solutions, we cannot measure or understand the effect of the complexity of these tasks on their performance. Further, external factors such as participant demographics may have affected the results of our study, though we have no evidence for this: upon performing a logistic regression with success as the dependent variable and the predictors being age, gender, professional role, years of experience and number of commits to the repository in the 4 months leading up to the study; none of the predictors were statistically significant.
Most participants in our study had not previously used Cursor or any other SWE agent, and thus our results mostly reflect early usage, rather than regular long-term use. Our study is also subject to observer effects; as participants knew that we were studying their interactions, they might may have used the tool differently than in a private setting. Likewise, our instruction that participants try to use AI for all their tasks means that, in the wild, developers would likely be more prone to abandonment. Variations in the duration of study sessions might also have affected participants’ behavior. Furthermore, while model call latency was the overall dominant latency bottleneck, our decision to have participants use a remote machine induced an additional UI latency which may have hindered their user experience.
Our choice of Cursor in early 2025 also has an effect on generalizability; while we found Cursor to have the most interaction features compared to other in-IDE SWE Agents and we anticipate our findings would apply to such tools, our findings may not generalize to non-IDE based SWE Agents. Further, since AI tools and models are constantly evolving, some findings reflect present-day limitations of agents and their underlying models.
@@ IV Results
Figure 3 visualizes participants’ success and time spent as they submitted 269 prompts to the agent across 33 issues. Participants were satisfied with the final patch for 16 of these tasks, while 4 of these tasks were abandoned for external reasons (such as participants being unconvinced of the issue’s validity) for a total success rate of 55% (16/(334))(16/(33-4)). In this section, we describe the strategies participants applied, the challenges they faced while collaborating with the agent, and finally how these strategies and other factors were associated with participants’ success.
@@ IV-A RQ1: How developers collaborate with SWE agents
In this subsection, we describe the patterns and strategies we observed in participants’ collaborations with the agent.
@@ IV-A1 Delegation Strategies
  • If the agent is unable to generate a comprehensive fix or verify its changes, participants must then manually both understand the generated fix and then iterate on it to get to a valid fix. We observed that this iteration can be time-consuming, as the it can entail not only understanding the code change itself, but also understanding the reasoning behind the change, navigating between code changes, and understanding changes to test files.
In the other approach (the incremental resolution strategy), participants manually divide the task into sequential sub-tasks and ask the agent to solve each sub-task sequentially in separate prompts. When the participant is satisfied with the solution to one sub-task, they then ask the agent to work on the next sub-task. Although this approach is safer than the one-shot strategy because mistakes made by the agent can be reviewed and caught earlier, it requires more proactive involvement and interaction from the participant to divide the issue into these sub-tasks. As evidence, participants who used this strategy used an average of 11.0 prompts per issue, compared to 7.0 for the one-shot strategy. Further, these participants were more likely to manually read existing code to enhance their own understanding while working on issues, doing so in 83% (15 out of 18) of the issues they worked on as compared to 60% (9 out of 15) for other participants.
One participant (P7), who had used Cursor before, explained that they preferred different strategies in different scenarios: “For an issue that is very contained, I trust it to come up with the changes itself, but for issues that are more involved with different parts of the codebase I feel have to point it in the right direction so that anything sensible comes out at the end”. However, we also observed that these approaches are not mutually exclusive; when delegating the entire issue to the agent via the one-shot strategy, an incomplete solution may require the participant to prompt the agent with extra steps via incremental resolution. As P3 mentioned, “The AI agent was able to generate good structure… once more context was needed, it committed errors that need more handholding and additional time to be spent on guiding it”.
@@ IV-A2 Providing Expert Information
We observed that participants provided two different types of information to the agent to help it make the desired code changes:
@@
  • Contextual information from the issue description and the environmental context (such as test logs). Eg: “There are build errors in your solution”
@@
  • Expert information relating to code implementation or convention that cannot be observed directly from context or the issue description and is based on the developer’s prior knowledge of the repository. Eg: “Usually we don’t set an upperbound. I think 3.10 should be fine now”
@@
Participants’ patterns of providing such information to the agent mirror those of software engineering managers. Kalliamvakou and colleagues [38] report that when assigning work, great managers in software engineering leave the implementation details to the engineer, providing actionable feedback when required. Analogously, participants tended to provide less expert information to the agent up-front, and more as feedback; they provided expert information in 49% of prompts that asked the agent to generate a new code change, compared to in 66% of prompts that ask the agent to refine a code change (i.e., iterate upon a previously generated change). This relationship was even stronger for the 12 participants in managerial roles – they provided expert insights in 45% of their prompts seeking new changes versus 75% when seeking to refine changes. As shown in Figure 4, the analogy is reflected in the post-study questionnaire, in which participants report viewing the agent more as an assistant (at a lower level in the organizational hierarchy) than as a colleague (at the same level) or advisor (at a higher level).
@@
While all participants usually tended to provide expert information in feedback prompts rather than up-front, we found that experience factors moderated how often participants provided such information. For example, participants who had more familiarity with the repository tended to provide expert information more often – those with over 10 commits to the repository in the four months leading up to the study provided such insights 70% of time when seeking any code changes as compared to 48% for other participants. Similarly, participants who had previous experience with Cursor also provided more expert insights, doing so in 81% of prompts seeking code changes. Further, participants applying the one-shot strategy provided expert insights less often (in 47% of their prompts seeking code changes) as compared to those applying the incremental resolution strategy (who did so in 64% of their prompts seeking code changes).
IV-A3 Requested Tasks
@@
We observed that participants not only requested the SWE agent to make code changes to resolve the task specified in the GitHub issue but also requested assistance in sub-tasks corresponding to other activities of the software development process. While 50% of participants’ prompts requested code changes, 27% of their prompts sought explanations of details related to the existing codebase (code comprehension), 16% of their prompts sought to run tests and 11% of their prompts sought explanations related to changes made by the agent (code review). Note that the above categories are not mutually exclusive. This distribution of prompts reflects participants’ perceptions of the agent as reported in the post-study questionnaire (Figure 4). Participants viewed the agent the most as a content generator; prompts requesting code generation were most common. Since ‘reference guide’ was a close second, participants also often used prompts to ask the agent about the codebase. On the other hand, participants did not view the agent as a reviewer, and were thus less likely to ask it to run tests or review its own code.
@@ IV-A4 Manual Actions
Participants’ perceptions of the agent shaped not only the types of prompts they provided the agent, but also the role the participants themselves took in the collaboration, tending to work on debugging and testing more than on writing code. Since participants reported not thinking of the agent as a code reviewer (Figure 4), they often took the responsibility to debug and test code upon themselves, doing so manually in 21 out of 33 issues by running tests or examining execution logs. In the case of bugs concerning UI changes, it was also easier for the participant to test changes manually by simply observing the state of the UI rather than uploading images to the chat for the agent to interpret.
On the other hand, since participants tended to view the agent more as a content generator, they were more likely to let the agent handle all the code changes to be made, manually writing code in only 14 issues. Further, in 10 of these issues, participants’ code interventions were limited to editing code suggested by the agent rather than adding new functionality.
Some participants expressed desire for AI tools to provide more support in non-code generation actions such as localizing, debugging, and testing, since these aspects of the software development process were often critical to resolving issues. P12 expressed that that the agent was not particularly helpful in localizing their issue, mentioning that “I think the main part [of solving the issue] was finding where to put the code, so it might have been marginally faster to do it myself”. Similarly, participants mentioned the challenges they faced when collaborating with the agent for debugging; for instance, P8 responded in the post-study questionnaire that the agent was “not too great or useful at the debugging and converting-a-repro-into-a-test loop”, and P11 responded that “It’s puzzling to me why [the agent is] so good at generating code but bad at dealing with basic environmental commands”. Participants’ manual actions thus reflected not only their perceptions of the agent, but also a forced response to the constrained capabilities of the agent as a collaborator.
IV-A5 Reviewing Agent Outputs
The agent provided three types of outputs to communicate with participants:
  • Execution: a trace of the agent’s work that appears while the agent is working, including details like files read and commands executed (the bulk of Figure 1).
@@ IV-A5 Reviewing Agent Outputs
We observed that participants tended to first track the agent’s execution live as-it-happened before reviewing explanations and diffs, with a preference for reading diffs. While Cursor Agent allows for users to synchronously perform manual tasks in parallel to its own execution, participants instead followed the agent’s execution as it happened in real time for 84% of their prompts. Usually, this live analysis was sufficient for participants to understand the agent’s process; after following the agent’s execution live, participants only reviewed the process that the agents followed (such as searching the codebase and reading files, in Figure 1b) for 7% of the agent’s responses. In contrast, for the majority of cases (specifically, after 75% of the agent’s responses), participants reviewed the agent’s outputs (diffs or explanations) after following its execution in real time.
When reviewing code outputs, participants preferred to verify diffs over change explanations. While participants followed the agent’s execution in real time for 95% of the responses that included code changes, these code changes often required further verification – participants reviewed diffs after 67% of such agent responses, while they only reviewed explanations after 31% of these responses. Only for 8% of responses with code changes did participants review code explanations without reviewing code diffs. This preference to review code directly is also reflected in the post-study questionnaire, in which participants ranked the ability to see detailed explanations of code changes as the least important feature among the different features of the agent (Figure 5). Manual verification of the agent’s outputs was usually sufficient for participants to have clarity on the agent’s code changes – they only asked for explanations of generated changes in 16% of their follow-up prompts.
@@
IV-A6 Iteration Patterns
Participants in our study often iterated on the agent’s outputs to get to a satisfactory solution, preferring to refine faulty patches rather than starting over. Iteration was required more often when working on feature requests – participants used an average of 11.2 prompts when working on feature requests as compared to 6.3 when working on bugs. Overall, participants used an average of 8.2 prompts per issue, out of which 52% of prompts that sought code changes sought refinements to previous changes by the agent, indicating that the agent’s first outputs were often unsatisfactory and required further modifications. Despite this, participants only rejected code changes 10% of the time after the agent generated code changes. Further, even though the agent allowed for participants to reset the chat to earlier points in the conversation (Figure 1a) – a feature we mentioned in the warm-up task – participants only used this feature in 15% of the issues they worked on. Additionally, participants only terminated the agent’s execution before completion (Figure 1d) in 11% of prompts. Thus, participants preferred to iterate on imperfect changes rather than starting afresh by stopping or resetting the agent.
Despite this proclivity to iterate on imperfect outputs, some participants indicated that they did not think that iterating was always a good strategy. P6 noted that the agent was “not understanding the implications of previously written bad code, and doubling down by adding more code instead of understanding the underlying issue.” Similarly, P14 mentioned of the agent in their post-study questionnaire that “if you continue to ask it to fix issues that it hallucinated on you might end up wasting time following in on the wrong path. It would be better to just explain the task in more detail or revert back to a previous breakpoint.”
IV-B RQ2: Barriers to Developer-Agent Collaborations
@@ IV-B1 Lack of tacit knowledge
The agent was sometimes an ineffective collaborator because it lacked tacit knowledge, which is gained through personal experience and is undocumented [5]. Unlike with expert insights related to implementation details, it may be difficult for participants to know when to discuss tacit knowledge, and even when they do, it may be difficult for them to articulate it in writing. Since the agent cannot understand non-verbal clues, this may hamper effective collaborations [40].
For example, in the case of P7 (who had prior experience with Cursor), the participant was able to interpret a temporal cue but unable to communicate their finding to the agent – when a terminal command running unit tests that was started by the agent seemed to be taking a long time to complete execution, the participant realized based on their experience with the repository that the tests were failing. However, rather than stopping the agent’s execution and explaining this in a prompt, they chose to work on the issue manually while the agent came to this realization itself. In other cases, the agent suffered from a lack of social context. For example, the agent once offered an explanation of the codebase that seemed to contradict comments on the issue by a maintainer of the repository, making participant P8 skeptical of the agent’s correctness. Prior knowledge of this context might have led the agent to question its assumptions before generating such an explanation.